# Sensing with submarine optical cables

### v3

@@ Abstract
In this paper, we establish the theoretical framework for understanding the sensing capabilities of optical fibers. We show the distinct advantage of polarization over phase in detecting subhertz environmental processes. Subsequently, we propose a scheme capable of extracting the spectrum of perturbations affecting a specific section at any position along an optical fiber by detecting the state of polarization of the backreflected light. We discuss two examples of earthquake detection and the detection of sea swells and ocean tides through the analysis of the state of polarization of an optical signal reconstructed by the receiver of a transoceanic cable, obtained from an online database.Zhan (2020) Finally, we provide the analytical expression for the cross-correlation of the polarization perturbations of two wavelength division multiplexed channels, and show that the analysis of the polarization correlations between adjacent channels can provide valuable insights into the localization of earthquakes.
@@ I Introduction
In this paper, we will first establish the theoretical basis for the understanding of the sensing capability of optical fibers. We will show the pros and cons of the use of the polarization-averaged optical phase and of the polarization as a sensing tool. We will propose a scheme that can provide the spectrum of the perturbation acting upon the cable by detecting the state of polarization of the backreflected light. We will then delve into the vast database in ref. Zhan, 2020 to show a few examples of earthquake detection using the reconstructed state of polarization of the receiver of the Curie cable system connecting Los Angeles, in California, to Valparaiso in Chile. Finally, we will present an analytical expression of the cross-correlation of the state of polarization between two wavelength division multiplexed channels and highlight how studying the correlations between the polarization of nearby channels may yield valuable insights into localization of earthquakes.
@@ II Basic equations
Let the tranverse electric field propagating in a single mode fiber be represented by the column vector
@@
representing an anticlockwise rotation by the angle $\theta(z)$ of in the xyx-y plane with respect to the $z$ axis, so that $\mathbf{V}(z)\,\mathbf{B}(z)\,\mathbf{V}^{\dagger}(z)$ produces a rotation by the same angle of the birefringence vector, and
@@
with the individual members of the product belonging to $\mathrm{SU}(2)$ (the special unitary group of degree 2, which is the group of unitary two-by-two matrices of unit determinant). Being $\mathrm{SU}(2)$ closed with respect to matrix multiplication, [^1] This property is an immediate consequence of the isomorphism between $\mathrm{SU}(2)$ and $\mathrm{SO}(3)$ (the group of three dimensional rotations around the origin of Stokes space) and of the fact that the concatenation of rotations is still a rotation. $\mathbf{U}(z)$ itself is a member of $\mathrm{SU}(2)$.Gordon and Kogelnik (2000); Sakurai and Napolitano (2017) Defining the polarization averaged phase as
$\overline{\varphi}_{0}=\frac{1}{2i}\,\mathrm{trace}\log\mathbf{U}_{0}=\frac{1}{2i}\log\det\mathbf{U}_{0},$ (20)
@@
so that the polarization averaged phase is the solution of Eq. (17). A member of the $\mathrm{SU}(2)$ group can always be expressed asGordon and Kogelnik (2000); Sakurai and Napolitano (2017)
@@
Being $\varphi_{0}(z)$ solution of Eq. (17), it is not affected by the rotation of the birefringence axes. Consequently, the fluctuations of the average phase are insensitive to the rotation of the birefringence axes, or equivalently, they are uncoupled to the fluctuations that leave unchanged the modulus of the birefringence eigenvalues. Consequence of this property is that, as we will see, fiber twist does not affect the polarization averaged phase. The evolution of the phase fluctuations can be expressed in terms of the polarization averaged wavevector by simply integrating Eq. (17)
@@
This vector describes in Stokes space the evolution of the field polarization. The equation describing the evolution of $\vec{s}(z)$ can be obtained using Eq. (13) in the definition of Stokes vector
$\vec{s}(z)=\langle\vec{s}(0)|\mathbf{U}^{\dagger}(z)\vec{\sigma}\mathbf{U}(z)|\vec{s}(0)\rangle$ (26)
differentiating both terms and using Eq. (18) and the equality $(\vec{a}\cdot\vec{\sigma})\vec{\sigma}=\vec{a}\mathds{1}+i\vec{a}\times\vec{\sigma}$ and that $\vec{\sigma}(\vec{a}\cdot\vec{\sigma})=\vec{a}\mathds{1}-i\vec{a}\times\vec{\sigma}$. The final result isGordon and Kogelnik (2000)
@@
Having derived the fundamental equations, the following section will be devoted to the analysis of the effect of external perturbations on the polarization averaged wavevector $\beta_{0}$ and on the fiber birefringence $\vec{\beta}$.
@@ III Wavevector perturbations
Let us consider a perfectly cylindrical fiber and treat any effects of the deviations from cylindrical symmetry (including birefringence) as a small perturbation. Assume first a perfectly cylindrical fiber with no preexisting birefringence. If some strain is applied to the fiber, the polarization averaged wavevector $\beta_{0}=2\pi n_{0}/\lambda$ turns into
@@
Here, $n_{0}$ is the unperturbed glass refractive index, $C$ is the photoelastic factor and $\epsilon(z)$ the strain. The strain also affects the propagation distance at every $z$, becoming $z(\epsilon)=[1+\epsilon(z)]z$. Thus, ${\operatorname{d}}z$ turns into ${\operatorname{d}}z^{\prime}(\epsilon)=[1+\epsilon(z^{\prime})]{\operatorname{d}}z^{\prime}$ and $\varphi_{0}$ changes to
@@
is the unperturbed wavevector. The effect of strain is to produce to the polarization averaged propagation constant the relative change
@@
We obtain the important result that in quasi-cylindrical symmetric fibers, in which birefringence is a small perturbation, the dependence of the birefringence on strain can be inferred from the dependence on strain of the polarization averaged wavevector, which is easier to characterize experimentally.Mecozzi et al. (2021)
@@ IV Phase-based sensing
requires that $|\Delta\varphi_{\mathrm{sig}}(T)|\gg\Delta\varphi_{\mathrm{rms}}$, where $\Delta\varphi_{\mathrm{rms}}$ is the root-mean square of the phase deviation due to the laser phase noise. The phase fluctuations of a quantum limited laser are described by a Wiener process with diffusion constant $D_{\varphi}=2\pi\nu$, where $\nu$ is the laser linewidth. After a time $T$ the root-mean square phase deviations originating from the laser phase noise are therefore
$\Delta\varphi_{\mathrm{rms}}(T)=\langle\Delta\varphi_{\nu}^{2}(T)\rangle^{1/2}=\sqrt{D_{\varphi}T}.$ (41)
The condition $|\Delta\varphi_{\mathrm{sig}}(T)|\gg\Delta\varphi_{\mathrm{rms}}$ becomes
$|\Delta\varphi_{\mathrm{sig}}(T)|\gg\sqrt{D_{\varphi}T}.$ (42)
If the rate of variation of the phase is $\dot{\varphi}$, then $\Delta\varphi_{\mathrm{sig}}(T)=\dot{\varphi}T$ and Eq. (42), after using the expression for $D_{\varphi}$ in terms of the linewidth, becomes
$|\dot{\varphi}|T\gg\sqrt{2\pi\nu T},$ (43)
@@
$\nu\ll\frac{\dot{\varphi}^{2}T}{2\pi}.$ (44)
A longer sampling time effectively averages the effects of laser phase noise, thereby reducing the requirements on the linewidth. However, the $2\pi$ periodicity of the phase imposes that $|\Delta\varphi_{\mathrm{sig}}|=|\dot{\varphi}|T\leq 2\pi$. Using for $T$ its maximum value $2\pi/|\dot{\varphi}|$, we obtain the somehow expected result
$\nu\ll|\dot{\varphi}|.$ (45)
This condition was met in the earthquake detection cases presented in ref. Marra et al., 2018, wherein the earthquakes induced a phase dynamics covering intervals on the order of thousands of radians over a timescale of tens of seconds, and the laser employed had sub-hertz equivalent linewidth. It is questionable whether the detection of processes with much weaker dynamics that occur over the same timescale or longer, like those associated with tsunami’s propagation, would be possible even utilizing lasers with similar ultra-low linewidth.
Let us now consider a different scenario, namely that the process is characterized by a bandwidth $1/(2T)$ and is recorded for a duration sufficient to utilize the short-time Fourier transform for its analysis.Oppenheim, Schafer, and Buck (1999) Let $\varphi_{0}(t)$ be the phase corrupted by the laser phase noise. The instantaneous frequency perturbed by the frequency noise of the laser has the expression
@@
We employed a normalization factor $1/n$ in the Fourier transform definition to ensure that the spectral amplitude $\tilde{\varphi}(\Omega_{\overline{k}})$ of a pure sinusoidal modulation at frequency $\Omega_{\overline{k}}$ is independent of the number of samples $n$. Fourier transforming both sides of Eq. (49) yields
@@
The conditions stated by Eqs. (58) and (59) are independent of the sampling time $T$ and the number of samples $n$ (provided that $n\gg 1$). Specifically, Eq. (59) illustrates that the maximum tolerable linewidth is directly proportional to the modulation frequency, the amplitude square of the phase modulation and the ratio $T_{\mathrm{win}}/T_{k}$, which is the number of temporal periods contained within the time window of the Fourier transform $T_{\mathrm{win}}=nT$. Therefore, widening the time window of the short-time Fourier transform $T_{\mathrm{win}}$ can alleviate the requirements on the laser linewidth compared to the detection of transient phase deviations discussed earlier in this section. However, this can only be achieved by sacrificing temporal resolution, which is determined by the amplitude of the time window $T_{\mathrm{win}}$. Consequently, the necessity of using ultra-stable lasers for detecting low intensity sub-hertz signals persists even with approaches based on the short-time Fourier transform.

### v4

@@ Abstract
In this paper, we establish the theoretical framework for understanding the sensing capabilities of megameter-long submarine optical cables. We show the distinct advantage of polarization over phase in detecting subhertz environmental processes. Subsequently, we propose a scheme capable of extracting the spectrum of perturbations affecting a specific section at any position along an optical fiber by detecting the state of polarization of the backreflected light. We discuss two examples of earthquake detection and the detection of sea swells and ocean tides through the analysis of the state of polarization of an optical signal reconstructed by the receiver of a transoceanic cable, obtained from an online database.Zhan (2020) Finally, we provide the analytical expression for the cross-correlation of the polarization perturbations of two wavelength division multiplexed channels, and show that the analysis of the polarization correlations between adjacent channels can provide valuable insights into the localization of earthquakes.
@@ I Introduction
Although optical sensing is a well-established field Lu et al. (2010); Pastor-Graells et al. (2016); Chen, Liu, and He (2017); Guerrier et al. (2020); Westbrook (2020) (see the excellent review provided by Ref. Lu et al., 2019 and references therein), in this paper we will mainly concentrate on techniques capable of providing sensing capabilities to submarine optical cables of megameter lengths. The paper is organized as follows. In Section II, we set up the basic equations describing the propagation of a polarized optical field in an unperturbed optical fiber. We introduce here the two quantities that play a central role for sensing, specifically the polarization-averaged phase and the polarization rotation vector, and show how these quantities are related to the fiber polarization-averaged wavevector and birefringence vector. In Sec. III, we consider the effect that time-dependent external perturbations produce on the polarization-averaged wavevector and on the birefringence vector of a fiber with random polarization coupling, like those universally used in optical communication systems. In Secs. IV and V, we set the theoretical basis for understanding how these perturbations can be used for sensing. In Sec. IV, we show the limitations imposed by the laser phase noise on the performance of phase-based sensing when applied to very low-frequency processes. We also show here the benefit of low-pass filtering the frequency deviations (and not the phase deviations) over the smallest possible bandwidth compatible with the signal bandwidth. In Sec. VI, we study how the signal accumulates along the link in sensing approaches based on transmission configurations, and show that while polarization is almost immune to the orientation of a seismic perturbation because the signal accumulation is incoherent, the phase signal is potentially sensitive to the orientation of a seismic wave, with higher sensitivity to waves orthogonal to the cable orientation in the perturbed area. This result may explain a trend observed in the outcomes of the experimental campaign reported in Ref. Donadello et al., 2024. Section VII is propedeutical to the analysis in Sec. VIII, and discusses the importance of using a rotating frame in the analysis of polarization perturbation. The use of this frame is equivalent to the use of the interaction representation in quantum mechanics.Sakurai and Napolitano (2017) In Sec. VIII, we propose a scheme that can extract the spectrum of the perturbation acting at a given position along the cable by detecting the state of polarization of the backreflected light, paving the way to the possibility of localizing the perturbation. Finally, in Sec. IX, we delve into the vast database in Ref. Zhan, 2020 to show a few examples of earthquake detection using the reconstructed state of polarization of the receiver of the Curie cable system connecting Los Angeles, California, to Valparaiso, Chile. In the same section, we also present an analytical expression of the cross-correlation of the state of polarization between two wavelength-division multiplexed channels and highlight how studying the correlations between the polarization of nearby channels may yield valuable insights into the localization of earthquakes.
@@ II Basic equations
Let the transverse electric field propagating in a single mode fiber be represented by the column vector
@@
representing an anticlockwise rotation by the angle $\theta(z)$ of in the $x-y$ plane with respect to the $z$ axis, so that $\mathbf{V}(z)\,\mathbf{B}(z)\,\mathbf{V}^{\dagger}(z)$ produces a rotation by the same angle of the birefringence vector, and
@@
with the individual members of the product belonging to $\mathrm{SU}(2)$ (the special unitary group of degree 2, which is the group of unitary two-by-two matrices of unit determinant). Being $\mathrm{SU}(2)$ closed with respect to matrix multiplication, [^1] This property is an immediate consequence of the isomorphism between $\mathrm{SU}(2)$ and $\mathrm{SO}(3)$ (the group of three dimensional rotations around the origin of Stokes space) and of the fact that the concatenation of rotations is still a rotation. $\mathbf{U}(z)$ itself is a member of $\mathrm{SU}(2)$.Gordon and Kogelnik (2000); Sakurai and Napolitano (2017) Performing the trace of $1/(2i)\log\mathbf{U}_{0}$ averages out the polarization dependent term of the argument of the exponential in $\mathbf{U}_{0}$ and is therefore equivalent to the average over all the polarization states.Gordon and Kogelnik (2000) Consequently, this operation returns the phase of the optical field averaged over all polarization states,
$\overline{\varphi}_{0}=\frac{1}{2i}\,\mathrm{trace}\log\mathbf{U}_{0}=\frac{1}{2i}\log\det\mathbf{U}_{0}.$ (20)
@@
Consequently, from now on we will refer to $\varphi_{0}$, defined as the solution of Eq. (17), as the polarization-averaged phase. A member of the $\mathrm{SU}(2)$ group can always be expressed asGordon and Kogelnik (2000); Sakurai and Napolitano (2017)
@@
Since the vector $\vec{\varphi}(z)$ contains all the information on the rotation of the field polarization from the input to the position $z$ along the fiber we will refer to this vector as the polarization rotation vector. Being $\varphi_{0}(z)$ solution of Eq. (17), it is not affected by the rotation of the birefringence axes. Consequently, the fluctuations of the average phase are insensitive to the rotation of the birefringence axes, or equivalently, they are uncoupled to the fluctuations that leave unchanged the modulus of the birefringence eigenvalues. Consequence of this property is that, as we will see, fiber twist does not affect the polarization-averaged phase. The evolution of the phase fluctuations can be expressed in terms of the polarization-averaged wavevector by simply integrating Eq. (17)
@@
This vector describes in Stokes space the evolution of the field polarization. The equation describing the evolution of $\vec{s}(z)$ can be obtained using Eq. (13) in the definition of Stokes vector and using that $\mathbf{U}_{0}(z)=\exp[i\varphi_{0}(z)]\mathbf{U}(z)$
$\vec{s}(z)=\langle\vec{s}(0)|\mathbf{U}^{\dagger}(z)\vec{\sigma}\mathbf{U}(z)|\vec{s}(0)\rangle,$ (26)
then differentiating both terms and using Eq. (18) and the equality $(\vec{a}\cdot\vec{\sigma})\vec{\sigma}=\vec{a}\mathds{1}+i\vec{a}\times\vec{\sigma}$ and that $\vec{\sigma}(\vec{a}\cdot\vec{\sigma})=\vec{a}\mathds{1}-i\vec{a}\times\vec{\sigma}$. The final result isGordon and Kogelnik (2000)
@@
Equation (24) shows that the value of the polarization-averaged phase is the effect of the accumulation of the infinitesimal contributions $\beta_{0}(z^{\prime}){\operatorname{d}}z^{\prime}$ and is therefore natural referring to $\beta_{0}(z^{\prime})$ as the polarization-averaged wavevector. Equation (27) in turn shows that the Stokes vector of the field evolves by accumulating infinitesimal rotations described by the rotation vectors $\vec{\beta}(z)$, and we will therefore refer to this vector as the birefringence vector. Having derived the fundamental equations, the following section will be devoted to the analysis of the effect of external perturbations on the polarization-averaged wavevector $\beta_{0}$ and on the birefringence vector $\vec{\beta}$.
@@ III Wavevector perturbations
Let us consider a perfectly cylindrical fiber and treat any effects of the deviations from cylindrical symmetry (including birefringence) as a small perturbation. Assume first a perfectly cylindrical fiber with no preexisting birefringence. If some strain is applied to the fiber, the polarization-averaged wavevector $\beta_{0}=2\pi n_{0}/\lambda$ turns into
@@
Here, $n_{0}$ is the unperturbed glass refractive index, $C$ is the photoelastic factor and $\epsilon(z)$ the strain. The strain also affects the propagation distance at every $z$, becoming $z(\epsilon)=[1+\epsilon(z)]z$. Thus, ${\operatorname{d}}z$ turns into ${\operatorname{d}}z^{\prime}(\epsilon)=[1+\epsilon(z^{\prime})]{\operatorname{d}}z^{\prime}$ and if we retain the symbol $\varphi_{0}$ for the polarization-averaged phase change in the absence of perturbation, and use the symbol $\varphi(z)$ for the value in the presence of perturbation, we get
@@
is the unperturbed wavevector. The effect of strain is to produce to the polarization-averaged propagation constant the relative change
@@
We obtain the important result that in quasi-cylindrical symmetric fibers, in which birefringence is a small perturbation, the dependence of the birefringence on strain can be inferred from the dependence on strain of the polarization-averaged wavevector, which is easier to characterize experimentally.Mecozzi et al. (2021)
@@ IV Phase-based sensing
requires that $|\Delta\varphi_{\mathrm{sig}}(T)|\gg\Delta\varphi_{\mathrm{rms}}$, where $\Delta\varphi_{\mathrm{rms}}$ is the root mean square of the phase deviation due to the laser phase noise. The phase fluctuations of a quantum limited laser are described by a Wiener process with diffusion constant $D_{\varphi}=2\pi\nu$, where $\nu$ is the laser linewidth, so that
$\langle\Delta\varphi_{\nu}^{2}(T)\rangle=D_{\varphi}T.$ (41)
Let us assume that we detect the phase over an observation time $T_{\mathrm{win}}=nT$ with sampling time $T$, and consider the instantaneous frequency deviations, which we define as the phase deviations divided by the time interval,
$\dot{\varphi}(T)=\frac{1}{T}\left[\Delta\varphi(z,t+T)-\Delta\varphi(z,t)\right].$ (42)
An estimate of the phase deviation over the time $T_{\mathrm{win}}$ can be obtained with the following procedure. We first take the average of the $n$ frequency estimates over the time interval. Being the signal and the additive phase noise of the laser independent, we may consider the effect of the averaging on the noise and on the signal separately. Let us consider the noise first. The root mean square of each term of the average is
$\dot{\varphi}_{\mathrm{rms}}(T)=\frac{1}{T}\sqrt{\langle\Delta\varphi_{\nu}^{2}(T)\rangle}=\sqrt{D_{\varphi}/T},$ (43)
and because of the independence of the increments of a Wiener process, the root mean square of the average over $n$ consecutive realizations is $(1/\sqrt{n})$ times the root mean square of each term, that is
$\dot{\varphi}_{\mathrm{rms}}(T,n)=\sqrt{D_{\varphi}/(nT)}=\sqrt{D_{\varphi}/T_{\mathrm{win}}}.$ (44)
This procedure is equivalent to applying a low-pass filter to the sequence with a bandwidth of the order of $1/T_{\mathrm{win}}$.
Let us now move to the analysis of the signal. If we assume that the rate of variation of the signal is constant over the observation time $T_{\mathrm{win}}$, the averaging does not affect the signal estimate and we have $\dot{\varphi}_{\mathrm{sig}}(T_{\mathrm{win}})=\dot{\varphi}$. The condition $|\dot{\varphi}_{\mathrm{sig}}(T_{\mathrm{win}})|\gg\dot{\varphi}_{\mathrm{rms}}(T_{\mathrm{win}})$ becomes, after using Eq. (44) and the expression of $D_{\varphi}$ in terms of the linewidth,
$|\dot{\varphi}|\gg\sqrt{2\pi\nu/T_{\mathrm{win}}},$ (45)
@@
$\nu\ll\frac{\dot{\varphi}^{2}T_{\mathrm{win}}}{2\pi}.$ (46)
Equation (46) shows the beneficial effect of a large observation time $T_{\mathrm{win}}$, which should be chosen to be as large as possible while remaining compatible with the bandwidth of the observed process, that is, with the maximum time over which $\dot{\varphi}$ can be assumed to be constant.
The analysis of Eq. (44) reveals that the root mean square of the frequency noise $\dot{\varphi}_{\mathrm{rms}}(T_{\mathrm{win}})$ depends only on $T_{\mathrm{win}}=nT$ and is independent of $n$ and $T$ separately. In particular, its value coincides with that for $n=1$, when no averaging is performed and $T_{\mathrm{win}}=T$ [^2] The property that the root mean square of the frequency noise is independent of the number of averaged samples $n$ is a direct consequence of the white noise character of the frequency noise of the laser. One may therefore be tempted to use a large sampling time and avoid filtering, using for the sampling time the desired value of $T_{\mathrm{win}}=T$. In this case, however, the $2\pi$ periodicity of the phase imposes that $|\Delta\varphi_{\mathrm{sig}}|=|\dot{\varphi}|T\leq 2\pi$. Using Eq. (46) with $T_{\mathrm{win}}=T$ and for $T$ its maximum value $2\pi/|\dot{\varphi}|$, we obtain
$\nu\ll|\dot{\varphi}|,$ (47)
which is a more restrictive condition than that given by Eq. (46).
It is very important to notice that inverting the order of the two operations of averaging (filtering) and differentiating makes the filtering ineffective. For a given $T_{\mathrm{win}}$, filtering of the frequency deviations does not reduce the frequency noise compared to the case without filtering and $T=T_{\mathrm{win}}$. Similarly, filtering of the phase noise does not reduce the phase fluctuations. The key benefit of low-pass filtering the frequency fluctuations is that it allows the use of longer integration times $T_{\mathrm{win}}$ (or equivalently narrower low-pass filters). Notice that, while the condition $|\dot{\varphi}|T\leq 2\pi$ must always be valid, the condition that the noise is smaller than the signal over the sampling time, $|\dot{\varphi}|T\gg\sqrt{2\pi\nu T}$, does not.
In principle, one could also avoid filtering following the phase over time intervals $T_{\mathrm{win}}$ that span many $2\pi$ periods using a sampling time sufficiently short to ensure that the phase changes do not exceed $2\pi$ over the sampling period. One can then calculate the frequency deviations by subtracting the initial values to the final values and dividing by $T_{\mathrm{win}}$. The condition for the linewith would still be Eq. (46). Besides this procedure being less reliable due to the need to unwrap the phase, it has the additional disadvantage of not filtering out other non-white noise components that might affect the detected frequency, originating either from the laser itself or from the detection process.
Our analysis shows the importance, in phase-based sensing that does not employ ultra-narrow linewidth lasers, of transforming the phase samples, obtained with a sampling time $T$ sufficiently small to avoid that the phase increments over $T$ exceed $2\pi$, into frequency samples, and the need for low-pass filtering the resulting frequency samples over a sufficiently narrow bandwidth. In practice, one should choose the narrowest possible bandwidth compatible with the bandwidth of the observed process.
Let us now check the condition (46) over the earthquake detection reported in Fig. 2 c) of Ref. Mazur et al., 2024. There, frequency oscillations of about 10 Hz amplitude can be clearly observed after a 60 mHz filtering of the phase traces. In Eq. (46), $1/T_{\mathrm{win}}$ is the cutoff frequency of the low-pass filter, hence a 60 mHz low-pass filtering is equivalent to averaging time $T_{\mathrm{win}}\simeq 1/0.06\simeq 17$ s. The condition for the laser linewidth Eq. (46) dictates that $\nu\ll 17\cdot 10^{2}/(2\pi)\simeq 265$ Hz, and this condition was satisfied by the linewidth of the fiber laser used in the study of Ref. Mazur et al., 2024, which was less than 100 Hz.
The above considerations apply to approaches based on the analysis of temporal traces. Let us now consider the case in which the phase is analyzed using instead a spectrogram, which is a time-frequency representation employing a short-time Fourier transform.Oppenheim, Schafer, and Buck (1999) Assume that the phase $\varphi_{0}(t)$ is characterized by a bandwidth $1/(2T)$. The instantaneous frequency perturbed by the frequency noise of the laser has the expression
@@
The normalization factor $1/n$ in the Fourier transform definition ensures that the spectral amplitude of a pure sinusoidal modulation $\varphi(t)=A\cos(\Omega_{\overline{k}}t)$ is $\tilde{\varphi}(\Omega_{\overline{k}})=(A/2)(\delta_{k,-\overline{k}}+\delta_{k,\overline{k}})$ and does not depend on the number of samples $n$. Fourier transforming both sides of Eq. (51) yields
@@
The conditions stated by Eqs. (60) and (61) are independent of the sampling time $T$ and the number of samples $n$, provided that $f_{k}\leq 1/(2T)$, which results from the condition $|\Omega_{k}|\leq|\Omega_{n/2}|$ in Eq. (55). Specifically, Eq. (61) illustrates that the maximum tolerable linewidth is directly proportional to the modulation frequency, the amplitude square of the phase modulation and the ratio $T_{\mathrm{win}}/T_{k}$, which is the number of temporal periods contained within the time window of the Fourier transform $T_{\mathrm{win}}=nT$. Therefore, widening the time window of the short-time Fourier transform $T_{\mathrm{win}}$ can alleviate the requirements on the laser linewidth compared to the detection of transient phase deviations discussed earlier in this section. However, this can only be achieved by sacrificing temporal resolution, which is determined by the amplitude of the time window $T_{\mathrm{win}}$. Consequently, the necessity of using ultra-stable lasers for detecting low intensity sub-hertz signals persists even with approaches based on the short-time Fourier transform.
@@
Equation (60) applies to all measurements using phase as a sensing probe, including distributed acoustic sensing. It does not however prevent the detection of millihertz signals if the oscillations have sufficient amplitude. For instance, Eq. (60) for the 1 mHz oscillations reported in Fig. 4 c) of Ref. Lindsey, Dawe, and Ajo-Franklin, 2019, obtained with $T_{\mathrm{win}}=7200$ s, yields $\nu\ll 0.045\,|\tilde{\varphi}|^{2}$. If the amplitude of the 1000 s period oscillations corresponds to a change of the optical path by 10 wavelengths, we have $|\tilde{\varphi}|\simeq 20\pi$, we obtain $\nu\ll 714$ Hz, which is a condition satisfied by the linewidth of good quality fiber lasers.Mazur et al. (2024)